# Beyond-Labels: Advancing Open-Vocabulary Segmentation With Vision-Language Models

### v3

@@ IV-A Dataset
We evaluated our approach on the PASCAL-5i [23] dataset, which is designed for few-shot segmentation approaches. It is based on the well-known PASCAL VOC 2012 dataset. The dataset is divided into four folds (i = 0, 1, 2, 3) and contains 20 different object classes. Each fold contained five classes, and the model was trained on the other fifteen classes. This cross-validation design separates the classes used for testing from the classes used for training, ensuring that this is indeed a valid evaluation strategy to determine how well few-shot learning algorithms generalize onto new categories. The support set contains some labeled instances, for example, 1-shot or 5-shot, from each of the five target classes in each fold to use as a reference for segmentation. Given this limited information in the support set, the query set should separate the objects precisely in the unlabeled photos that contain objects from the same five classes. Models trained on 15 classes and tested on the remaining 5 are doing so to cover all classes iteratively, using a technique called ”leave-one-fold-out.” Since this is indeed an exhaustive evaluation framework, it has made PASCAL-5i a standard benchmark for the comparison of few-shot segmentation. We have reported the results of our method in the ablation study and results comparison section.
@@ IV-B Evaluation Metric
Our key evaluation metric is mean intersection-over-union (mIoU), which averages IoU across all classes within the fold.
@@ IV-C Results Comparison
Table 2 compares different models on the segmentation task for all folds in terms of the overall mean IoU, namely, mIoU. SPNet has a representation with an overall mean IoU of 18.3 and folds 2 and 3 performed the weakest, showing the low ability of the network to generalize well against the unseen complexities of classes within few-shot segmentation tasks. ZS3Net performed better with a mIoU of 38.3, trying to generalize across folds in semantic features The OSLSM improves further by achieving a mIoU of 40.8, which is superior to that of ZS3Net and SPNet; however, challenges remain for particular folds such as 3. Our model surpasses all baselines by a mIoU of 41.5, with maximum performance at folds 0 and 1 at 44.6 and 49.4, respectively. This again supports the hypothesis that integrating Fourier embedding with multi-model fusion, together with a decoder, will improve feature representation by paying full consideration to the mutual image and textual information. This indicates that the relative improvement of 3.2 in mIoU over OSLSM and 8.4 over ZS3Net demonstrates the effectiveness of our proposed components in handling complex spatial relationships and unseen class distributions. These results confirm that our approach not only achieves better segmentation accuracy but also ensures more consistent performance across diverse folds.
@@ IV-D Ablation Study
We investigated the contribution of each core component in the proposed ”Beyond-Labels” framework. Specifically, we investigated the impact of the Fourier Embedding (F_Emb), multi-model embedding fusion module (Fusion), and the use of a decoder instead of traditional positional embedding techniques. Detailed results of this study can be seen in Table 1 for breaking down performance in different configurations.
@@
  • By default, it uses only an upward model that includes the usual positional embedding and upsampling using a decoder. The use of this standard baseline model resulted in minimum performance, by which B_L_0 obtained a mIoU of only 25.7. The large decreases in scores in folds 2 and 3 represent a limitation based on the use of standard position embedding without robust feature fusion in action. Without Fourier embedding, the model cannot capture such high-dimensional and expressive positional information, which is critical for improving segmentation accuracy, especially in challenging folds such as folds 2 and 3.
@@
  • While for B_L_1, incorporation of Fourier Embedding improves its performance on all counts with the mIoU increasing to 36.9. This, in turn, indicates that adding frequency-based representation is highly vital for boosting the feature extraction required for improving the generalization of non-visible classes. More important is the model’s capability of adaptation to different complicated spatial scenarios due to improved performances on all folds.
@@
  • The best-performing model, B_L_2, is the full model, including the Fourier embedding, Fusion, and decoder, with a mIoU of 41.5. Additional positional information is introduced by Fourier Embedding that, combined with the fusion module, enables the model to learn richer representations. Further refinement of the upsampling process with the decoder leads to consistent improvements in all the folds. This includes the Fourier embedding module with the Fusion module, which generates high-dimensional and expressive forms of position-based information, thus yielding significant gains in segmentation accuracy. This further facilitates better spatial understanding within the image, boosting the capability for more precise feature representations of the model.
@@
The benefits of Fourier embedding are easily found in all folds, where traditional embeddings often fall short. This can be integrated into the full model to demonstrate how higher-quality positional encoding can be combined with other components to achieve top-level performance.
IV-E Implementation Details
The vision-language model (CLIP [20]) is frozen throughout the training process. Each of the 8 levels that make up the multi-model embedding fusion module has 4 attention heads. After k convolutional layers, the decoder performs a 2× upsampling operation. We employ the AdamW optimizer with a cosine scheduler to decay the learning rate to maximize training. We use cross-entropy for the loss function and set the temperature factor at ($\mathcal{\tau}$)= 0.08.
IV-F Conclusion
@@
Our research aims to make advanced vision and language models more helpful for open-vocabulary semantic segmentation. We introduce ”Beyond-Labels,” a compact and efficient module that links image and language models to do open-vocabulary semantic segmentation, which involves finding and categorizing items in images, even for previously unknown categories. We tested it on the “PASCAL” dataset and found that it performs extremely well, exceeding some existing approaches. Despite its simplicity, “Beyond-Labels” delivers excellent results.
IV-G Future Work
To assess our work’s resilience and enhance generalization across many domains, we will expand it to include more datasets and perform more ablation studies. Through testing on various datasets, we hope to ensure that our method works reliably and successfully in practical situations, capturing a wider range of difficulties and subtleties in semantic segmentation tasks.
IV-H Template List

### v5

@@ IV-A Dataset
We evaluate our approach on the PASCAL-5i [21] dataset, which is designed for few-shot segmentation approaches. It is based on the well-known PASCAL VOC 2012 [9] and SDB [12] datasets. The dataset is divided into four folds (i = 0, 1, 2, 3) and contains 20 different object classes. Each fold contained five classes, and the model is trained on the other fifteen classes. This cross-validation design separates the classes used for testing from the classes used for training, ensuring that this is indeed a valid evaluation strategy to determine how well few-shot learning algorithms generalize to new categories. The support set contains some labeled instances, for example, 1-shot or 5-shot, from each of the five target classes in each fold to use as a reference for segmentation. Given this limited information in the support set, the query set should separate the objects precisely in the unlabeled photos that contain objects from the same five classes. Models trained on 15 classes and tested on the remaining 5 are doing so to cover all classes iteratively, using a technique called ”leave-one-fold-out.” Since this is indeed an exhaustive evaluation framework, it has made PASCAL-5i a standard benchmark for the comparison of few-shot segmentation. We have reported the results of our method in the ablation study and the results comparison section.
@@ IV-B Evaluation Metric
The key evaluation metric is mean intersection-over-union (mIoU), which averages IoU across all classes within the fold.
@@ IV-C Results Comparison
In terms of mIoU, Table 1 indicates a comparison of different models of segmentation on all folds. The worst scores belonged to Folds 2 and 3, indicating the incapability of SPNet to go against the native complexities of classes in few-shot segmentation tasks based on generalization power, where the overall mIoU reached 18.3. ZS3Net outperformed SPNet with a score of 38.3 on mIoU for generalizing semantic features across folds. OSLSM outperforms ZS3Net and SPNet with an mIoU of 40.8; however, there are problems with some folds, such as fold 3. OSLSM also compared their method’s performance against a range of baselines, including LogReg, Siamese, Finetuning, and 1-NN, and adopted previous work on dense pixel prediction as baselines against which they compare. PANet and CO-FCN are both competitive, with PANet having an mIoU of 49.1 and CO-FCN having an mIoU of 41.1, outperforming OSLSM and ZS3Net. Improvements are also seen in AMP-2 and SG-One, with AMP-2 obtaining an mIoU of 43.4 and SG-One getting 46.3 and outperforming in multiple folds (folds 1 and 2). With an mIoU of 49.4, our model outperforms all baselines; it performs best at folds 0 and 2, respectively. This again supports the assumption that integrating Fourier embedding with multimodal fusion, along with a visual decoder, improved feature representation by providing maximum consideration to the reciprocal textual and visual information. This indicates that the increase of 8.6 in mIoU compared to OSLSM, 5.3 compared to PANet, and 11.1 compared to ZS3Net demonstrates the effectiveness of our proposed components in handling complex spatial relationships and unseen class distributions. These results confirm that not only does our approach achieve more precise segmentation, but it also achieves more balanced performance across folds.
@@ IV-D Ablation Study
We investigated the contribution of each core component in the proposed ”Beyond-Labels” framework. Specifically, we investigated the impact of the Fourier Embedding (F_Emb), multimodal embedding fusion module (Fusion), and the use of a decoder instead of traditional positional embedding techniques. Detailed results of this study can be seen in Table 3 for breaking down performance in different configurations.
@@
  • By default, it uses only an upward model that includes the usual positional embedding and upsampling using a decoder. Using this standard baseline model resulted in minimum performance, by which B_L_0 obtained an mIoU of 34.0. The large decreases in scores in fold 3 represent a limitation based on the use of standard position embedding without robust feature fusion in action. Without Fourier embedding, the model cannot capture high-dimensional and expressive positional information, critical for improving segmentation accuracy, especially in challenging folds such as folds 2 and 3.
@@
  • While for B_L_1, incorporation of Fourier embedding improves its performance on all counts, with the mIoU increasing to 39.5. This, in turn, indicates that adding frequency-based representation is vital for boosting the feature extraction required to improve the generalization of non-visible classes. More important is the model’s adaptation capability to different complicated spatial scenarios due to improved performance on all folds.
@@
  • The best-performing model, B_L_2, is the full model, including the Fourier embedding, fusion, and decoder, with an mIoU of 49.4. Additional positional information is introduced by Fourier Embedding that, combined with the fusion module, enables the model to learn richer representations. Further refinement of the upsampling process with the decoder leads to consistent improvements in all the folds. This includes the Fourier embedding module with the Fusion module, which generates high-dimensional and expressive forms of position-based information, thus yielding significant gains in segmentation accuracy. This further facilitates better spatial understanding within the image, boosting the capability for more precise feature representations of the model.
@@
The benefits of Fourier embedding are easily found in all folds in B_L_1 and B_L_2, where traditional embeddings often fall short.
V Conclusion
@@
Our research aims to make advanced vision and language models more helpful for open-vocabulary semantic segmentation. We introduce ”Beyond-Labels,” a compact and efficient module that links vision and language models to do open-vocabulary semantic segmentation, which involves finding and categorizing items in images, even for previously unknown categories. We tested it on the “PASCAL” dataset and found that it performs extremely well, exceeding some existing approaches. Despite its simplicity, “Beyond-Labels” delivers excellent results. To assess our work’s resilience and enhance generalization across many domains, we will expand it to include more datasets and perform more ablation studies. Through testing on various datasets, we hope to ensure that our method works reliably and successfully in practical situations, capturing a wider range of difficulties and subtleties in semantic segmentation tasks.
VI Appendix
VI-A Template List